# Challenging Forgets: Unveiling the Worst-Case Forget Sets in Machine Unlearning

### v1

@@ 3 Preliminaries, Motivation, and Problem Statement
Figure 2: Performance of the MU methods Retrain and $\ell_{1}$-sparse under random and worst-case forgetting scenarios at different forgetting data ratios. Variance over 10 random forget set selections is indicated by the shaded areas of the dashed lines.
Fig. 2 provides a motivating example showcasing the sensitivity of MU evaluation to the (random) forget set selection. We examine two unlearning approaches: Retrain (exact unlearning) and $\ell_{1}$-sparse (a state-of-the-art approximate unlearning method), and explore two evaluation scenarios: traditional random data forgetting and our proposed worst-case forget set evaluation, which will be elaborated in Sec. 4. For a given unlearning method and forget set selection, we assess the unlearning effectiveness and model utility of the unlearned model ${{\bm{\theta}}_{\mathrm{u}}}$ following the evaluation metrics used in [8]. The effectiveness metrics include unlearning accuracy (UA), calculated as 1–the accuracy of ${{\bm{\theta}}_{\mathrm{u}}}$ on the forget set $\mathcal{D_{\mathrm{f}}}$, and MIA (membership inference attack) on $\mathcal{D_{\mathrm{f}}}$, which determines whether a data point in $\mathcal{D_{\mathrm{f}}}$ was used in training ${{\bm{\theta}}_{\mathrm{u}}}$. The utility metrics include remaining accuracy (RA), measured by the accuracy of ${{\bm{\theta}}_{\mathrm{u}}}$ on the retain set $\mathcal{D_{\mathrm{r}}}$, and testing accuracy (TA), the accuracy of ${{\bm{\theta}}_{\mathrm{u}}}$ on the test set.
@@ 5.2 Experiment Results
Validating the worst-case forget set via Retrain. We begin by justifying the worst-case unlearning performance of the chosen forget set through the exact unlearning method, Retrain. In Table 2, we examine the performance disparities between the worst-case forget set and the random forget set in the task of MU for image classification on CIFAR-10, when employing Retrain at different forgetting data ratios including $1\%$, $5\%$, $10\%$, and $20\%$. In terms of unlearning effectiveness, the chosen worst-case forget set consistently poses the greatest challenge for unlearning in all scenarios tested, as indicated by a significant drop in UA and MIA to nearly 0% (see the ‘Worst-case’ and ‘Diff’ columns of Table 2). In addition, the variance in worst-case unlearning effectiveness performance (as measured by UA and MIA) remains significantly lower than that observed with random data forgetting at various forgetting data ratios. Furthermore, the utility of the unlearned model, as indicated by RA and TA, shows no loss when comparing unlearning on worst-case forget sets to random forget sets. Intriguingly, the TA of models unlearned with the worst-case forget set may even surpass those unlearned with random sets, hinting at a connection to coreset selection that will be further explored later. In Appendix C, we broaden our evaluation of the worst-case forget set across additional datasets and model architectures. The findings align with those presented in Table 2.
@@
Analyzing the selected worst-case forget set through a coreset lens. Our prior observation from Table 2 indicates that TA of models unlearned with worst-case forget sets could surpass those subjected to random forget sets. This prompts us to investigate the characteristics of data typically selected within the worst-case forget sets. In what follows, we explore this question through the lens of coreset, given by a subset of the training data deemed sufficient for model training [76, 58, 55]. We hypothesize that the data points included in the worst-case forget set likely do not constitute the coreset, as forgetting them poses challenges, possibly owing to their strong inherent correlation with data not selected for forgetting. Motivated by this hypothesis, we explore whether the complement of the worst-case forget set corresponds to the coreset.
To answer the above question, Fig. 3 contrasts the testing accuracy of an image classifier (ResNet-18) trained on the complement of the worst-case forget set (termed ‘Worst’) against that trained on coresets identified through the moderate coreset selection method (termed ‘Moderate’) [57]. For comparison, we also present the performance using the model trained on the original full dataset (termed ‘Original’) and the random set (termed ‘Random’). As we can see, for both CIFAR-10 and CIFAR-100 datasets, training models on the complement of the worst-case forget set achieve superior TA across various data selection ratios, outperforming both random data selection and the coreset selection method, Moderate. Furthermore, TA achieved by using the complement of the worst-case forget set on CIFAR-100 can even exceed the performance of the original model trained on the entire dataset at $90\%$ and $95\%$ selection ratios (i.e., $10\%$ and $5\%$ forgetting ratios). These observations suggest that the chosen worst-case forget set indeed does not constitute a coreset, whereas its complement serves as a coreset.
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
An extended study: Worst-case forget set on prompt-wise unlearning. Extending from data-wise forgetting, we further demonstrate the efficacy of our approach in prompt-wise forgetting for text-to-image generation. We utilize the stable diffusion (SD) model [75] on the UnlearnCanvas dataset, a benchmark dataset designed to evaluate the unlearning of painting styles and objects [12]. In UnlearnCanvas, a text prompt used as the condition of image generation is given by ‘A painting of [Object Name] in [Style Name] Style.’ We considered 10 objects and 10 styles (100 combinations in total) for prompts and selected 10% of them to unlearn. For designated prompts targeted for unlearning, we apply the Erased Stable Diffusion (ESD) [15] technique; See Appendix B for further implementation details. Fig. 5 presents examples of images generated using the pre-trained SD model, the unlearned model by forgetting a random prompt set, and the unlearned model by forgetting the identified worst-case forget set. For each model, images are generated under two conditions, given from (1) the unlearned prompt set $P_{\text{u}}^{(\text{w/r})}$ where ${}^{(\mathrm{w})}$ or ${}^{(\mathrm{r})}$ indicates the worst-case or the random forget set, and (2) the normal prompt set $P_{\text{n}}$ irrelevant to forget sets. As we can see, the unlearned diffusion model is unable to prevent image generation based on prompts from the worst-case forget set ($P_{\text{u}}^{(\text{w})}$), resulting in similar image outputs to those of the original diffusion model. In contrast, the diffusion model unlearned through random forgetting can avoid generating accurate images based on the unlearning-targeted prompts ($P_{\text{u}}^{(\text{r})}$) from the random forget set, displaying a significant deviation from the original model’s outputs, indicating successful unlearning. Furthermore, when conditioned on the normal, forgetting-irrelevant prompts ($P_{\text{n}}$), both worst-case and random forgetting-oriented diffusion models perform well, generating the requested images. The above results indicate that erasing the influence of prompts from the worst-case forget set introduces new challenges of MU for image generation. We refer readers to Appendix C for more visualizations.
Additional results. In Appendix C, we also demonstrate the effectiveness of worst-case forget set in the scenario of class-wise forgetting on ImageNet [74], and expand our study to the easiest-case unlearning.
6 Conclusion
@@
In this study, we delved into the challenge of pinpointing the worst-case forget set in MU (machine unlearning), introducing a fresh perspective that broadens the scope and enhances the effectiveness of MU beyond conventional methods like random data forgetting. By employing BLO (bi-level optimization), we developed a structured approach to accurately identify these pivotal sets. Utilizing sign-based gradient unrolling, we show that the BLO-focused task of determining the worst-case forget set is solvable through first-order optimization techniques, with applicability extending to class-wise and prompt-wise forgetting scenarios. Through extensive experiments, we demonstrated the effectiveness of our proposed method in different data-model setups, showcasing its significance for improved reliability in MU evaluations. In addition, our case studies in biased dataset unlearning and prompt-wise forgetting for generative models highlight the adaptability and depth of our approach. This work also paves the way for future inquiries into machine unlearning, advocating for a shift towards more comprehensive and robust evaluation paradigms.
@@ A.1 The closed-form projection operation
According to [78, Proposition 1], the solution of the above projection problem yields
@@ C.4 Transferability of worst-case forget sets between different models and methods
Figure A2: UA of various unlearning methods on CIFAR-10 using ResNet-18 with a forgetting ratio of 10%. The rightmost column represents UA on random forget set, while other columns depict UA on worst-case forget set.
In this section, we validate the transferability of worst-case forget sets across a wider range of model architectures and methods.
Concerning the transferability between models, we leverage a diverse range of models for the selection process, including ResNet-18, ResNet-50 [71], VGG-16, and VGG-19 [73]. Conversely, for the evaluation process, we employ various models and adopt Retrain as the corresponding unlearning method. The UA (unlearning accuracy) results are exhibited in Fig A1. Notably, when the worst-case forget set is selected using one model and subsequently undergoes unlearning with another model, the UA shows significantly lower than that of random forget set. This observation clearly demonstrates the transferability of worst-case forget set across diverse models.
Regarding the transferability between methods, we employ various approximate unlearning objectives in the selection process to specify the lower-level optimization problem (9), while utilizing different unlearning methods during evaluation. The UA results are illustrated in Fig A2. As evident from the figure, the columns corresponding to the four methods, FT, $\ell_{1}$-sparse, RL, and SalUn, exhibit deeper shades than the column for random, indicating lower UA values. Consequently, FT, $\ell_{1}$-sparse, RL, and SalUn are more suitable for addressing the lower-level problem.
@@ C.5 From worst-case unlearning to easiest-case unlearning
C.6 Identifying worst-case forget set in class-wise forgetting.
Table A5: Performance of various MU methods on ImageNet, ResNet-18. The content format follows Table 3.

### v2

@@ 3 Preliminaries, Motivation, and Problem Statement
Figure 2: Performance of the MU methods Retrain and $\ell_{1}$-sparse under random and worst-case forgetting scenarios at different forgetting data ratios on CIFAR-10 using ResNet-18. Variance over 10 random forget set selections is indicated by the shaded areas of the dashed lines.
Fig. 2 provides a motivating example showcasing the sensitivity of MU evaluation to the (random) forget set selection. We examine two unlearning approaches: Retrain (exact unlearning) and $\ell_{1}$-sparse (a state-of-the-art approximate unlearning method), and explore two evaluation scenarios: traditional random data forgetting and our proposed worst-case forget set evaluation, which will be elaborated in Sec. 4. For a given unlearning method and forget set selection, we assess the unlearning effectiveness and model utility of the unlearned model ${{\bm{\theta}}_{\mathrm{u}}}$ following the evaluation metrics used in [8]. The effectiveness metrics include unlearning accuracy (UA), calculated as 1–the accuracy of ${{\bm{\theta}}_{\mathrm{u}}}$ on the forget set $\mathcal{D_{\mathrm{f}}}$, and MIA (membership inference attack) on $\mathcal{D_{\mathrm{f}}}$, which determines whether a data point in $\mathcal{D_{\mathrm{f}}}$ was unused in training ${{\bm{\theta}}_{\mathrm{u}}}$. The utility metrics include remaining accuracy (RA), measured by the accuracy of ${{\bm{\theta}}_{\mathrm{u}}}$ on the retain set $\mathcal{D_{\mathrm{r}}}$, and testing accuracy (TA), the accuracy of ${{\bm{\theta}}_{\mathrm{u}}}$ on the test set.
@@ 5.2 Experiment Results
Validating the worst-case forget set via Retrain. We begin by justifying the worst-case unlearning performance of the chosen forget set through the exact unlearning method, Retrain. In Table 2, we examine the performance disparities between the worst-case forget set and the random forget set in the task of MU for image classification on CIFAR-10, when employing Retrain at different forgetting data ratios including $1\%$, $5\%$, $10\%$, and $20\%$. In terms of unlearning effectiveness, the chosen worst-case forget set consistently poses the greatest challenge for unlearning in all scenarios tested, as indicated by a significant drop in UA and MIA to nearly 0% (see the ‘Worst-case’ and ‘Diff’ columns of Table 2). In addition, the variance in worst-case unlearning effectiveness performance (as measured by UA and MIA) remains significantly lower than that observed with random data forgetting at various forgetting data ratios. Furthermore, the utility of the unlearned model, as indicated by RA and TA, shows no loss when comparing unlearning on worst-case forget sets to random forget sets. Intriguingly, the TA of models unlearned with the worst-case forget set may even surpass those unlearned with random sets, hinting at a connection to coreset selection that will be further explored later. In Appendix C.4, we broaden our evaluation of the worst-case forget set across additional datasets and model architectures. The findings align with those presented in Table 2.
@@
Analyzing the selected worst-case forget set through a coreset lens. Our prior observation from Table 2 indicates that TA of models unlearned with worst-case forget sets could surpass those subjected to random forget sets. This prompts us to investigate the characteristics of data typically selected within the worst-case forget sets. In what follows, we explore this question through the lens of coreset, given by a subset of the training data deemed sufficient for model training [76, 58, 55]. We hypothesize that the data points included in the worst-case forget set likely do not constitute the coreset, as forgetting them poses challenges, possibly owing to their strong inherent correlation with data not selected for forgetting. Motivated by this hypothesis, we explore whether the complement of the worst-case forget set corresponds to the coreset. Fig. 3 contrasts the testing accuracy of an image classifier (ResNet-18) trained on the complement of the worst-case forget set (termed ‘Worst’) against that trained on coresets identified through other coreset selection methods, including Moderate [57], EL2N [52] and GraNd [52]. For comparison, we also present the performance using the model trained on the original full dataset (termed ‘Original’). As we can see, for both CIFAR-10 and CIFAR-100 datasets, training models on the complement of the worst-case forget set achieves TA comparable to or even better than the state-of-the-art coreset selection methods across various data selection ratios. Furthermore, TA achieved by using the complement of the worst-case forget set on CIFAR-100 can even exceed the performance of the original model trained on the entire dataset at $90\%$ and $95\%$ selection ratios (i.e., $10\%$ and $5\%$ forgetting ratios). These observations suggest that the chosen worst-case forget set indeed does not constitute a coreset, whereas its complement serves as a coreset.
@@
Refer to captionRefer to caption Refer to captionRefer to caption
@@
Refer to captionRefer to caption Refer to captionRefer to caption
@@
Refer to captionRefer to caption Refer to captionRefer to caption
@@
Refer to captionRefer to caption Refer to captionRefer to caption
@@
An extended study: Worst-case forget set on prompt-wise unlearning. Extending from data-wise forgetting, we further demonstrate the efficacy of our approach in prompt-wise forgetting for text-to-image generation. We utilize the stable diffusion (SD) model [75] on the UnlearnCanvas dataset, a benchmark dataset designed to evaluate the unlearning of painting styles and objects [12]. In UnlearnCanvas, a text prompt used as the condition of image generation is given by ‘A painting of [Object Name] in [Style Name] Style.’ We considered 10 objects and 10 styles (100 combinations in total) for prompts and selected 10% of them to unlearn. For designated prompts targeted for unlearning, we apply the Erased Stable Diffusion (ESD) [15] technique; See Appendix B for further implementation details. Fig. 5 presents examples of images generated using the pre-trained SD model, the unlearned model by forgetting a random prompt set, and the unlearned model by forgetting the identified worst-case forget set. For each model, images are generated under two conditions, given from (1) the unlearned prompt set $P_{\text{u}}^{(\text{w/r})}$ where (w) or (r) indicates the worst-case or the random forget set, and (2) the normal prompt set $P_{\text{n}}$ irrelevant to forget sets. As we can see, the unlearned diffusion model is unable to prevent image generation based on prompts from the worst-case forget set ($P_{\text{u}}^{(\text{w})}$), resulting in similar image outputs to those of the original diffusion model. In contrast, the diffusion model unlearned through random forgetting can avoid generating accurate images based on the unlearning-targeted prompts ($P_{\text{u}}^{(\text{r})}$) from the random forget set, displaying a significant deviation from the original model’s outputs, indicating successful unlearning. Furthermore, when conditioned on the normal, forgetting-irrelevant prompts ($P_{\text{n}}$), both worst-case and random forgetting-oriented diffusion models perform well, generating the requested images. The above results indicate that erasing the influence of prompts from the worst-case forget set introduces new challenges of MU for image generation. We refer readers to Appendix C for more visualizations.
Additional results. In Appendix C, we demonstrate the effectiveness of worst-case forget set in the scenario of class-wise forgetting on ImageNet [74]. We also demonstrate the generality of the identified worst-case forget set across different models and unlearning methods. Furthermore, we examine the uniqueness of the worst-case forget set by mixing it with other randomly selected data points for unlearning, and show that the proposed method can be easily adapted to address the easiest-case unlearning as well.
6 Conclusion and Discussion
@@
In this study, we delved into the challenge of pinpointing the worst-case forget set in MU (machine unlearning), introducing a fresh perspective that broadens the scope and enhances the effectiveness of MU beyond conventional methods like random data forgetting. By employing BLO (bi-level optimization), we developed a structured approach to accurately identify these pivotal sets. Utilizing sign-based gradient unrolling, we show that the BLO-focused task of determining the worst-case forget set is solvable through first-order optimization techniques, with applicability extending to class-wise and prompt-wise forgetting scenarios. Through extensive experiments, we demonstrated the effectiveness of our proposed method in different data-model setups, showcasing its significance for improved reliability in MU evaluations. In addition, our case studies in biased dataset unlearning and prompt-wise forgetting for generative models highlight the adaptability and depth of our approach.
Although our worst-case performance assessment was inspired by the lack of robustness in random data forgetting, it also deepens the understanding of when MU becomes ‘easy’ or ‘difficult’ and the underlying reasons from a data selection-based interpretability perspective. Additionally, our results encourage rethinking the role of data difficulty in unlearning. For example, incorporating a curriculum based on these difficulty levels may significantly impact unlearning performance. We term the incorporation of curriculum learning [78] into machine unlearning as curriculum unlearning, which may show promise in improving unlearning effectiveness. Furthermore, the inability of retraining from scratch to unlearn the identified challenging forget set prompts a reevaluation of its appropriateness in defining ‘exact’ unlearning.
@@ A.1 The closed-form projection operation
According to [79, Proposition 1], the solution of the above projection problem yields
@@ C.4 Transferability of worst-case forget sets between different models and methods
Figure A2: UA of unlearning methods on CIFAR-10 using ResNet-18 with a 10% forgetting ratio. The rightmost column shows UA on random forget set; others show UA on worst-case forget set.
Table A4: Performance of various unlearning methods on CIFAR-10 using ResNet-18 with a 10% forgetting ratio under worst-case forget sets obtained using RL. For performance under random forget sets, refer to Table 3. The result format follows Table 3.
Methods Worst-Case Forget Set
UA RA TA MIA Avg. Gap
Retrain 0.02 100.00 94.67 0.16 0
Relabeling-free
FT 0.44 (0.42) 92.94 (7.06) 87.93 (6.74) 0.44 (0.29) 3.63
EU-$k$ 1.36 (1.33) 97.99 (2.01) 90.75 (3.92) 11.93 (11.78) 4.76
SCRUB 5.04 (5.02) 89.92 (10.08) 85.70 (8.97) 20.71 (20.56) 11.16
$\ell_{1}$-sparse 0.04 (0.02) 97.59 (2.41) 91.62 (3.05) 0.13 (0.02) 1.38
Relabeling-based
RL 4.62 (4.60) 99.96 (0.04) 93.59 (1.08) 98.36 (98.20) 25.98
BE 38.38 (38.36) 77.00 (23.00) 71.42 (23.25) 96.64 (96.49) 45.27
BS 31.02 (31.00) 81.88 (18.12) 75.62 (19.05) 96.82 (96.67) 41.21
SalUn 2.09 (2.07) 99.98 (0.02) 94.00 (0.67) 97.40 (97.24) 25.00
In this section, we validate the transferability of worst-case forget sets across a wider range of model architectures and methods. Concerning the transferability between models, we leverage a diverse range of models for the selection process, including ResNet-18, ResNet-50 [71], VGG-16, and VGG-19 [73]. Conversely, for the evaluation process, we employ various models and adopt Retrain as the corresponding unlearning method. The UA (unlearning accuracy) results are exhibited in Fig. A1. Notably, when the worst-case forget set is selected using one model and subsequently undergoes unlearning with another model, the UA shows significantly lower than that of random forget set. This observation clearly demonstrates the transferability of worst-case forget set across diverse models.
Regarding the transferability between methods, we employ various approximate unlearning objectives in the selection process to specify the lower-level optimization problem (9), while utilizing different unlearning methods during evaluation. The UA results are illustrated in Fig. A2. As evident from the figure, the columns corresponding to the four methods, FT, $\ell_{1}$-sparse, RL, and SalUn, exhibit deeper shades than the column for random, indicating lower UA values. Consequently, FT, $\ell_{1}$-sparse, RL, and SalUn are more suitable for addressing the lower-level problem. In Table A4, we further test the unlearning methods under the worst-case forget set obtained using RL to perform the lower-level unlearning process in BLO. The results are consistent with Fig. A2.
@@ C.5 From worst-case unlearning to easiest-case unlearning
Refer to caption Refer to caption
(a) Uniqueness of worst-case forget set (b) Mixture of worst-case forget set
Figure A3: UA of Retrain on CIFAR-10 using ResNet-18. (a) UA under worst-case forgetting scenarios at different forgetting data ratios. (b) UA under mixture of random and worst-case forgetting scenarios at different mixture ratios.
C.6 Uniqueness and mixture of worst-case forget set
To verify the uniqueness of worst-case forget set, we identified worst-case forget set for different forgetting data ratios and performed unlearning using Retrain. We found that a maximal set with zero UA can exist. As shown in Fig. A3-(a), with appropriately defined set sizes (up to $34\%$ of the entire dataset), our method consistently identifies a worst-case forget set with 00 UA.
Furthermore, any subset of this set will also exhibit the worst-case property. Fig. A3-(b) illustrates that including any part of the worst-case set complicates the unlearning process. When the forget set represents a $34\%$ ratio comprising a mix of worst-case forget set and random forget set and unlearning is performed using Retrain, the unlearning becomes increasingly difficult as the proportion of worst-case random forget set increases, which is indicated by the decrease in UA. This highlights the importance and significance of worst-case forgetting.
C.7 Identifying worst-case forget set in class-wise forgetting.
Table A6: Performance of various MU methods on ImageNet, ResNet-18. The content format follows Table 3. ($\bullet$) after Retrain in Worst-Case Forget Set indicates the difference from Random Forget Set.